OpenStudy (anonymous):

Two semi trucks are driving loads from Chicago to Denver, distance of 1125 miles. First truck leaves at 8 AM, averages 50 mph. Second truck leaves at 9 AM, averages at 54 mph. How much would the driver of the first truck have to increase her speed in order to arrive in Denver first?

OpenStudy (anonymous):

the answer is "at least 1.53 mph" . But I don't know how to get to this.

OpenStudy (phi):

You have to think about what is going on. (It helps to have already solved similar problems) First truck leaves at 8 AM, averages 50 mph. After one hour, the truck is 50 miles closer to Denver, so it has 1125 miles - 50 miles= 1075 miles to go

OpenStudy (phi):

Second truck leaves at 9 AM, averages at 54 mph, and has 1125 miles to go Just to get an idea of what is going on, I would figure out how long will it take the first truck to get to Denver, starting at 9 AM. use rate*time= distance, or time = distance/rate time = 1075/50= 21.5 hours the second truck will take 1125/54 = 20.83 hours so that says the second truck will get to Denver first. unless the first truck speeds up

OpenStudy (phi):

If we are the first truck, we say, "we are only allowed 20 and 5/6 hour to go 1075 miles" now use rate* time = distance : rate * 125/6 = 1075 (btw, 20.8333... and 125/6 and 20 and 5/6 are all the same number) solve for rate r \[ r \cdot \frac{125}{6} = 1075 \] can you do that ?

OpenStudy (anonymous):

51.6?

OpenStudy (phi):

yes, that is what I got. the 1st truck was going 50, and it has to speed up to 51.6 so it must increase its speed by 1.6 mph I don't think the answer can be 1.53 mph ??

OpenStudy (phi):

or, maybe they are thinking what should the new speed be starting immediately? i.e. at 8 AM the 1st truck has to get to denver in 20 5/6 hours ? can you figure that out ?

OpenStudy (anonymous):

I'll try. Thank you, though!

OpenStudy (phi):

wait!, because the 1st truck starts 1 hour earlier, it has 21 and 5/6 hours

OpenStudy (phi):

so the problem is distance = 1125 miles rate is unknown time is 21 and 5/6 hours = 131/6 hours rate * 131/6 = 1125